Method of Substitution
x+y-z=-1
8x-6y+4z=32
2x-2y-33z=5

Found 2 solutions by Edwin McCravy, TigerAlgebra:
Answer by Edwin McCravy(20055)

 x +  y -   z = -1
8x - 6y +  4z = 32
2x - 2y - 33z =  5

Surely you didn't copy this correctly, because
the solution to this system is

x = 452/241, y = -662/241,  z = 31/241

No teacher in their right mind would assign a
problem with such a ridiculous solution as that.

Edwin

Answer by TigerAlgebra(40)

{x,y,z} = {452/241,-662/241,31/241}
See the complete solution for system of linear equations x+y-z=-1
8x-6y+4z=32
2x-2y-33z=5